A 6% semiannual coupon bond matures in 4 years. The bond has a face value of $1,000 and a current yield of 6.5996%. What are the bond's price and YTM? (Hint: Refer to Footnote 6 for the definition of the current yield and to Table 7.1) Do not round intermediate calculations. Round your answer for the bond's price to the nearest cent and for YTM to two decimal places.

Bond’s price: $

YTM:
%
Answer #1

Annual Coupon=6%*1000 =60
Bond Price =Annual Coupon/Current Yield =60/6.5996% =909.1460

Semi annual Coupon =60/2 =30
Number of Periods =4*2 =8
YTM using financial calculator
N=8;PMT =30;PV =-909.1460;FV =1000;CPT I/Y =4.37%

YTM =2*4.37% =8.74%